# NVM-in-Cache: Repurposing Commodity 6T SRAM Cache into NVM Analog Processing-in-Memory Engine using a Novel Compute-on-Powerline Scheme

### v1

@@ III-C PIM Mode: Compute on Powerlines
Let us assume the RRAM devices have been programmed to the desired state corresponding to the weights of a deep learning network. The PIM phase of operation corresponds to the in-memory dot-product computation, which is executed in two cycles, each lasting 3.5 ns. The core idea is to exploit the programmed resistance states of the RRAM devices as weights and the applied wordline values as input activations (IA), while the resulting output current is collected along the power lines (VDD1/VDD2). Since VDD1 and VDD2 are shared across a column of the cache array, the output currents from multiple cells naturally accumulate in the current domain, thereby enabling efficient MAC operations. In the first cycle, WL1 and WL2 are held at 0 V for the initial 1.5 ns, while BL is driven to 0.8 V. During this period, the gated-GND control signals V1 and V2 are maintained at 0.8 V. Simultaneously, VDD1 is switched from 0.8 V to the reference voltage provided by the weighted configuration circuit (WCC), which performs the weighted current addition, while VDD2 remains at 0.8 V. For the subsequent 1 ns, WL1 is driven by the IA, V1 and V2 are pulled to 0, and the output current is measured at the VDD1 line. This is followed by 1 ns during which VDD1 and V1 are restored to the nominal value of 0.8 V. Finally, V2 is also returned to 0.8 V, restoring SRAM hold. The second cycle follows a similar sequence. WL1 and WL2 are set to 0 V for the first 1.5 ns, and BLB is driven to 0.8 V. V1 and V2 remain at 0.8 V during this period, while VDD2 is switched from 0.8 V to the bias voltage provided by the WCC, with VDD1 held constant at 0.8 V. In the next 1 ns, WL2 is applied with the IA, V1 and V2 are pulled to 0, and the output current is sampled at the VDD2 line. This is followed by 1 ns during which VDD2 and V2 are restored to 0.8 V. Finally, V1 is also returned to 0.8 V, completing the cycle and re-establishing SRAM hold. The output current through VDD1 and VDD2 are sampled only when IA is applied at WL1 and WL2, thereby performing analog dot product operation.
@@
Consider the case where the stored values at the internal nodes are Q = 1 and QB = 0, as shown in Fig. 5(a). Under this condition, the PMOS transistor M2 is switched ON, while M4 remains OFF. During the first 1.5 ns, VDD1 is pulled to a reference voltage provided by the WCC, which simultaneously discharges Q from 1 (0.8 V) to the reference voltage while keeping WL1 = 0 and VDD2 = 0.8 V. This interval of 1.5 ns is used as a precaution since VDD1 is shared across 128 cells along a column, and the parasitic capacitances require time to settle. In the subsequent 1 ns, if WL = 1 (IA = 1), the programmed resistance state of $R_{LEFT}$ determines the current magnitude: either a large current (LRS) or a small current (HRS) flows through VDD1. With V2 disabled (0 V), QB is held at 0 V, thereby ensuring data retention during computation. In the following 1 ns, both V1 and VDD1 are restored to 0.8 V. Since M2 remains ON (as QB = 0), Q recharges to 0.8 V, completing the cycle. During the second cycle, for the first 1.5 ns, Q remains at logic 1, which keeps M4 OFF and QB at 0. In the next 1 ns, WL2 = 1 and BLB = 1, charging QB to logic 1. In this case, NMOS transistor M3 turns ON, but since V1 = 0 V, Q is not discharged. With M4 OFF, only minimal current is observed at VDD2. In the following 1 ns, when V2 is restored to 0.8 V, NMOS M5 switches ON and discharges QB, thereby preserving the original data. After completing the computation cycle, the circuit transitions back to the SRAM hold state.
@@
For IA = 0, during the first 1.5 ns, Q discharges from 0.8 V to the reference voltage. Consequently, during the sampling window (when IA = 0 is applied to WL1), there is no voltage difference between Q and VDD1, resulting in negligible current flow through the power lines and ensuring a dot-product result of 0. Thus, when Q = 1 and QB = 0, the dot-product output is obtained on the left-hand side, whereas when Q = 0 and QB = 1, the output is obtained on the right-hand side. In this way, the two-cycle operation enables dot-product computation regardless of the data stored within the SRAM cell. During the in-memory dot product operation, WL1 and WL2 are activated for 1 ns (first and second cycle). Under this condition, if IA = 1, a direct current path from the bitline (BL) to ground is established, resulting in excess power consumption. To address this issue, V1 and V2 are activated only after the wordlines are deasserted, during the subsequent 1 ns interval. This approach allows the bit-cell to alternate seamlessly between conventional SRAM functionality and in-memory MAC operations, while preserving data integrity.
@@ IV-B Multi-bit PIM Computing
To enable multi-bit computation within the proposed 6T-2R cache-based architecture, we implement a $4\text{b}\times 4\text{b}$ MAC operation between the stored weights and IA for a 128-row activation simultaneously. Each 4-bit word corresponds to the weight storage, while the input activation is shared across the rows. The computation is carried out using a bit-serial scheme, which requires four cycles to complete the multi-bit operation. Although a bit-parallel approach using a digital-to-analog converter (DAC) could in principle execute the multi-bit IA in a single cycle, such an implementation would introduce significant peripheral overhead in conventional cache architectures. Specifically, it would also increase the design complexity required for accurate quantization of the accumulated current from the MAC array. For this reason, we adopt the bit-serial approach as a more area- and power-efficient strategy [32].
@@ IV-C Mapping of Weights and Input Activations
We adopt the IFM reuse mapping strategy described in [33]. In the proposed 6T–2R architecture, the IA are applied along the rows (wordlines, WL1 and WL2), while the outputs are also collected across the columns (VDD1 and VDD2). The weights are stored in the 6T–2R array, as illustrated in Fig. 7, while the IFMs are mapped column-wise according to the kernel dimensions. Specifically, $K\times K$ submatrices are generated, and the corresponding input data for each kernel position is directed to its associated submatrix. Each submatrix is mapped onto a 128 $\times$ 128 6T–2R bank (sub-array). During the first cycle, IFMs are multiplied with the 3D kernels mapped across the banks, producing partial sums that are accumulated outside the array to generate the initial OFMs. In the following cycle, the IFMs shift by one stride. Importantly, IFMs from the first cycle are reused in the second cycle as neighboring banks forward the required IFMs. Repeating this process over multiple cycles produces the final OFM.
@@ V-A Bit-cell and Word Layout
The layout study is conducted using the GlobalFoundries 22 nm FDSOI PDK. The standard 6T SRAM cell occupies an area of around 0.1 um2 [34], with wordlines routed separately, VSS routed horizontally, and VDD routed vertically. Within the bit-cell, the RRAM devices are symmetrically placed on top of the PMOS source terminals and the VDD lines. A 4-bit word is formed by connecting four such SRAM bit-cells, where the VSS lines are shared row-wise and subsequently connected to the gated GND transistors (V1 and V2), as illustrated in Fig. 8.
@@
Figure 9: Comparison of conventional 6T SRAM and proposed 6T-2R bitcell: (a) RRAM I–V characteristics, (b) Hold SNM, (c) Read SNM, and (d) Write SNM. The proposed design shows comparable stability.
@@
We next examine the robustness of the proposed design in terms of static noise margins (SNM). Fig. 9 (b)-(c) presents a comparative analysis of read, write, and hold SNMs for the proposed 6T-2R bit-cell against a conventional 6T SRAM. Despite the addition of two RRAM elements and the gated GND control transistors, the proposed design exhibits only marginal changes in the noise margins compared to the baseline 6T cell. During the hold operation, the butterfly curve shows that the proposed bit-cell retains nearly identical static noise margins as the conventional design, thereby confirming that the integration of non-volatile elements does not compromise the bistability of the cross-coupled inverters. In the read mode, the proposed cell demonstrates a slight reduction in SNM compared to the 6T SRAM, due to the additional series resistance introduced by the RRAM devices. However, this degradation remains minimal and well within acceptable limits for reliable operation. Although the write SNM is somewhat reduced compared to the 6T SRAM, this trade-off is acceptable, and writability can be further enhanced through conventional write-assist techniques. The read latency of the baseline 6T SRAM is $660~\text{ps}$, while the proposed 6T-2R incurs a slightly higher latency of $686~\text{ps}$, reflecting minimal impact of the altered bitcell. Similarly, the read energy for a $512$-bit row increases modestly from $2.23~\text{fJ}$ in the conventional 6T SRAM to $3.34~\text{fJ}$ in the proposed design. This small overhead is a reasonable trade-off for the additional non-volatile and in-memory computing functionality.
Figure 10: Linearity characteristics of the proposed 6T-2R bit-cell across process corners for 128 row activation: (a) weight vs. accumulated voltage, and (b) weight vs. sampled voltage, for $S$, $T$, and $F$ corners.
Figure 11: Linearity characteristics of the proposed 6T-2R bit-cell (a) across process corners for 128 row activation: weight vs. accumulated current; (b) Change in current vs number of rows activated.
Figure 12: ADC characterization of the proposed 6T-2R PIM for 128 row activation: (a) weight-to-ADC output transfer curves with calibrated ($V_{\text{REFP}}=660$ mV, $V_{\text{REFN}}=90$ mV) and uncalibrated ($V_{\text{REF}}=800$ mV) SAR ADC, and (b) ADC output response as a function of accumulated MAC value.
@@ V-C Linearity Analysis
The linearity of the proposed 6T-2R bit-cell was evaluated by sweeping the programmed weight (w=0w=01515) and recording both the accumulated column current and the corresponding sense voltage across different process corners. Fig. 10 shows the weight-to-voltage characteristics before and after sampling. This showcase that sample and hold circuit does not contribute to any non-linearity. Fig. 11(a) showcase the current linearity at various corners. In the TT corner, the current increases nearly linearly. A similar trend is observed at the SS corner. At the FF corner exhibits a deviation from linearity. This nonlinearity is attributed to the stronger transistor drive in the FF corner, which reduces the effective voltage swing across the RRAM stack and limits incremental current contributions, but this nonlinearity can be accounted for while training the neural network model. Monotonicity is preserved across all corners, ensuring correct weight ordering. These observations emphasize the importance of calibration to maintain accuracy under fast-corner conditions. Fig. 11 (b) illustrates the change in current with the number of row activations.
Figure 13: Variation of output: (a) Voltage; (b) Current for 128-row activation during PIM mode.
The analog output of the proposed $128\times 512$ array was digitized using a 6-bit SAR ADC operating at $50~\text{MHz}$. Fig. 12(a) shows the weight-to-digital transfer characteristics for both calibrated and uncalibrated cases. In the uncalibrated mode, the ADC output spans only codes 7–48 across the 6-bit output range, thereby utilizing less than 70% of the available dynamic range. This compression not only reduces the effective resolution but also introduces a systematic offset that distorts the weight-to-code mapping. After calibration, the reference voltages in the CDAC as shown in Fig. 6(d) were tuned to $V_{\text{REFP}}=820~\text{mV}$ and $V_{\text{REFN}}=260~\text{mV}$. With this optimization, the full 6-bit code space (0–63) is exercised across the 4-bit weight range. As shown in Fig. 12(b) after post-processing, the calibrated response eliminates the offset present in the uncalibrated case and restores a near-ideal linear mapping. Each weight increment corresponds to an average step of $\sim 4$ ADC codes, thereby preserving monotonicity and maximizing quantization efficiency. Fig. 13 shows the variation in output voltage and current when 128 rows are activated simultaneously for a change in 1 LSB in Monte Carlo simulations. These results highlight that reference tuning and calibration are essential to fully exploit the ADC’s dynamic range when interfaced with the analog outputs of the proposed 6T-2R array.
@@ V-D Throughput and Energy Efficiency
The proposed 6T-2R subarray consists of $128\times 512$ bit-cells and follows the input-activation and weight mapping strategy described in the previous section, with both input activation and weight precision set to 4 bits. The system latency is dominated by the SAR ADC, which operates at 50 MHz and incurs a delay of 160 ns. Due to the bit-serial computation approach, the latency for performing PIM using $R_{LEFT}$ is 640 ns, and the same applies for $R_{RIGHT}$. After completing two PIM computations, the normalized throughput achieved is 0.4 TOPS at 0.8 V. The overall system energy is primarily dominated by the 6T-2R array, which accounts for approximately 60% of the total energy consumption, followed by the ADC and the WCC block. The normalized energy efficiency of the 6T-2R array is calculated to be 491.78 TOPS/W. In terms of area, the total macro size is approximately 0.1 mm2, with the ADC occupying nearly 70% of the area. The resulting normalized compute density is estimated to be 4.37 TOPS/mm2.
@@
Parameter TCASII’24[35] ISSCC’23[36] ISSCC’22[37] TCASI’23[38] TCASI’23[39] JSSCC’24[40] This Work
Technology
180nm CMOS 28nm FDSOI 22nm FDSOI 65nm CMOS 28nm FDSOI 22nm FDSOI 22nm FDSOI
Array Size
8Kb 16Kb 256Kb 101Kb 16Kb 256Kb 64Kb
PIM Domain
Time Charge Current Charge Charge Current Current
Memory Type
6T SRAM
+ 9T
10T1C
SRAM
1T1R
RRAM
10T1C
SRAM
6T
SRAM
1T1R
MRAM
6T-2R
SRAM
+RRAM
Cache Data
Retention
No No No No No No Yes
Dataset
CIFAR-10 NA CIFAR-10 CIFAR-10 CIFAR-10 CIFAR-10 CIFAR-10
Accuracy
86.1% NA 91.74% 88.6% 85.07% 90.25% 91.27%
Input/Weight
Precision
8/8 8/8 8/8 8/8 4/4 4/4 4/4
Output Precision
@@
Throughput
0.07 7.65 142.2 12.8 12.8 54.3 25.6
Energy Efficiency
(TOPS/W)
0.291 16.02 0.96 10.3 16.1 5.26 30.73
Normalized
Throughput
@@
[…]
Normalized
Energy Efficiency
[…]
Normalized
Compute Density
@@
Baseline (no ADC nonlinearity or noise) 91.84
ADC nonlinearity only (fine-tuned) 91.55
ADC nonlinearity + noise (fine-tuned) 91.27

### v2

@@ III-C PIM Mode: Compute on Powerlines
Let us assume the RRAM devices have been programmed to the desired state corresponding to the weights of a deep learning network. The PIM phase of operation corresponds to the in-memory dot-product computation, which is executed in two cycles, each lasting 3.5 ns. The core idea is to exploit the programmed resistance states of the RRAM devices as weights and the applied wordline values as input activations (IA), while the resulting output current is collected along the power lines (VDD1/VDD2). Since VDD1 and VDD2 are shared across a column of the cache array, the output currents from multiple cells naturally accumulate in the current domain, thereby enabling efficient MAC operations.
In the first cycle, WL1 and WL2 are held at 0 V for the initial 1.5 ns, while BL is driven to 0.8 V. During this period, the gated-GND control signals V1 and V2 are maintained at 0.8 V. Simultaneously, VDD1 is switched from 0.8 V to the reference voltage provided by the weighted configuration circuit (WCC), which performs the weighted current addition, while VDD2 remains at 0.8 V. For the subsequent 1 ns, WL1 is driven by the IA, V1 and V2 are pulled to 0, and the output current is measured at the VDD1 line. This is followed by 1 ns during which VDD1 and V1 are restored to the nominal value of 0.8 V. Finally, V2 is also returned to 0.8 V, restoring SRAM hold. The second cycle follows a similar sequence. WL1 and WL2 are set to 0 V for the first 1.5 ns, and BLB is driven to 0.8 V. V1 and V2 remain at 0.8 V during this period, while VDD2 is switched from 0.8 V to the bias voltage provided by the WCC, with VDD1 held constant at 0.8 V. In the next 1 ns, WL2 is applied with the IA, V1 and V2 are pulled to 0, and the output current is sampled at the VDD2 line. This is followed by 1 ns during which VDD2 and V2 are restored to 0.8 V. Finally, V1 is also returned to 0.8 V, completing the cycle and re-establishing SRAM hold. The output current through VDD1 and VDD2 are sampled only when IA is applied at WL1 and WL2, thereby performing analog dot product operation.
@@
Consider the case where the stored values at the internal nodes are Q = 1 and QB = 0, as shown in Fig. 5. Under this condition, the PMOS transistor M2 is switched ON, while M4 remains OFF. During the first 1.5 ns (starting at 8 ns in Fig. 5(b)), VDD1 is pulled to a reference voltage provided by the WCC, which simultaneously discharges Q from 1 (0.8 V) to the reference voltage while keeping WL1 = 0 and VDD2 = 0.8 V. This interval of 1.5 ns is used as a precaution since VDD1 is shared across 128 cells along a column, and the parasitic capacitances require time to settle. In the subsequent 1 ns, if WL = 1 (IA = 1), the programmed resistance state of $R_{LEFT}$ determines the current magnitude: either a large current (LRS) or a small current (HRS) flows through VDD1. With V2 disabled (0 V), QB is held at 0 V, thereby ensuring data retention during computation. In the following 1 ns, both V1 and VDD1 are restored to 0.8 V. Since M2 remains ON (as QB = 0), Q recharges to 0.8 V, completing the cycle. During the second cycle, for the first 1.5 ns, Q remains at logic 1, which keeps M4 OFF and QB at 0. In the next 1 ns, WL2 = 1 and BLB = 1, charging QB to logic 1. In this case, NMOS transistor M3 turns ON, but since V1 = 0 V, Q is not discharged via M3. With M4 OFF, only minimal current is observed at VDD2. In the following 1 ns, when V2 is restored to 0.8 V, NMOS M5 switches ON and discharges QB, thereby preserving the original data. After completing the computation cycle, the circuit transitions back to the SRAM hold state.
@@
For IA = 0, during the first 1.5 ns, Q discharges from 0.8 V to the reference voltage. Consequently, during the sampling window (when IA = 0 is applied to WL1), there is no voltage difference between Q and VDD1, resulting in negligible current flow through the power lines and ensuring a dot-product result of 0. Thus, when Q = 1 and QB = 0, the dot-product output is obtained on the left-hand side, whereas when Q = 0 and QB = 1, the output is obtained on the right-hand side. In this way, the two-cycle operation enables dot-product computation regardless of the data stored within the SRAM cell. This approach allows the bit-cell to alternate seamlessly between conventional SRAM functionality and in-memory MAC operations, while preserving data integrity.
@@ IV-B Multi-bit PIM Computing
To enable multi-bit computation within the proposed 6T-2R cache-based architecture, we implement a $4\text{b}\times 4\text{b}$ MAC operation between the stored weights and IA for a 128-row activation simultaneously. Each 4-bit word corresponds to the weight storage, while the input activation is shared across the rows. The computation is carried out using a bit-serial scheme, which requires four cycles to complete the multi-bit operation. Although a bit-parallel approach using a digital-to-analog converter (DAC) could in principle execute the multi-bit IA in a single cycle, such an implementation would introduce significant peripheral overhead in conventional cache architectures. Specifically, it would also increase the design complexity required for accurate quantization of the accumulated current from the MAC array. For this reason, we adopt the bit-serial approach as a more area- and power-efficient strategy [bitserial].
@@ IV-C Mapping of Weights and Input Activations
We adopt the IFM reuse mapping strategy described in [mappingiscas]. In the proposed 6T–2R architecture, the IA are applied along the rows (wordlines, WL1 and WL2), while the outputs are also collected across the columns (VDD1 and VDD2). The weights are stored in the 6T–2R array, as illustrated in Fig. 7, while the IFMs are mapped column-wise according to the kernel dimensions. Specifically, $K\times K$ submatrices are generated, and the corresponding input data for each kernel position is directed to its associated submatrix. Each submatrix is mapped onto a 128 $\times$ 128 6T–2R bank (sub-array). During the first cycle, IFMs are multiplied with the 3D kernels mapped across the banks, producing partial sums that are accumulated outside the array to generate the initial OFMs. In the following cycle, the IFMs shift by one stride. Importantly, IFMs from the first cycle are reused in the second cycle as neighboring banks forward the required IFMs. Repeating this process over multiple cycles produces the final OFM.
@@ V-A Bit-cell and Word Layout
The layout study is conducted using the GlobalFoundries 22 nm FDSOI PDK. The standard 6T SRAM cell occupies an area of around 0.1 um2 [gfsram], with wordlines routed separately, VSS routed horizontally, and VDD routed vertically. Within the bit-cell, the RRAM devices are symmetrically placed on top of the PMOS source terminals and the VDD lines. A 4-bit word is formed by connecting four such SRAM bit-cells, where the VSS lines are shared row-wise and subsequently connected to the gated GND transistors (V1 and V2), as illustrated in Fig. 8.
@@
Figure 9: Comparison of conventional 6T SRAM and proposed 6T-2R bitcell: (a)-(b) Write Hold, (c)-(d) Read, and (e)-(f) Hold Static Noise Margin (SNM).
@@
We next examine the robustness of the proposed design in terms of static noise margins (SNM). Fig. 9 presents a comparative analysis of read, write, and hold SNMs for the proposed 6T-2R bit-cell for different resistance state against a conventional 6T SRAM. Despite the addition of two RRAM elements and the gated GND control transistors, the proposed design exhibits only marginal changes in the noise margins compared to the baseline 6T cell. During the hold operation, the butterfly curve shows that the proposed bit-cell retains nearly identical static noise margins as the conventional design, thereby confirming that the integration of non-volatile elements does not compromise the bistability of the cross-coupled inverters. In the read mode, the proposed cell demonstrates a slight reduction in SNM compared to the 6T SRAM, due to the additional series resistance introduced by the RRAM devices. However, this degradation remains minimal and well within acceptable limits for reliable operation which could be enhanced by read-assist technique. Although the write SNM is somewhat reduced compared to the 6T SRAM, this trade-off is acceptable, and writability can be further enhanced through conventional write-assist techniques. The read latency of the baseline 6T SRAM is $660~\text{ps}$, while the proposed 6T-2R incurs a slightly higher latency of $686~\text{ps}$, reflecting minimal impact of the altered bit-cell. This small overhead is a reasonable trade-off for the additional non-volatile and in-memory computing functionality.
Figure 10: Linearity characteristics of the proposed 6T-2R bit-cell across process corners for 128 row activation: (a) weight vs. accumulated output voltage and current, and (b) normalized output current vs number of row activations.
Figure 11: ADC characterization of the proposed 6T-2R PIM for 128 row activation: (a) weight-to-ADC output transfer curves with calibrated ($V_{\text{REFP}}=570$ mV, $V_{\text{REFN}}=155$ mV) and uncalibrated ($V_{\text{REF}}=800$ mV) SAR ADC, and (b) ADC output response as a function of accumulated MAC value.
@@ V-C Linearity Analysis
The linearity of the proposed 6T-2R bit-cell was evaluated by sweeping the programmed weight (w=0w=01515) and recording both the accumulated column current and the corresponding sense voltage. Fig. 10(a) shows the weight-to-voltage characteristics after sampling. Fig. 10(b) showcase the normalized output current output vs number of row activation. The amount of nonlinearity is attributed to the stronger transistor drive at different corners, which reduces the effective voltage swing across the RRAM stack and limits incremental current contributions, but this nonlinearity can be accounted for while training the neural network model. Monotonicity is preserved across all corners, ensuring correct weight ordering.
Figure 12: Variation of change in output: (a) Voltage; (b) Current for 128-row activation with weight of 15 during PIM mode.
The analog output of the proposed $128\times 512$ array was digitized using a 6-bit SAR ADC operating at $50~\text{MHz}$. Fig. 11(a) shows the weight-to-digital transfer characteristics for both calibrated and uncalibrated cases. In the uncalibrated mode, the ADC output spans only codes 7–48 across the 6-bit output range, thereby utilizing less than 70% of the available dynamic range. This compression not only reduces the effective resolution but also introduces a systematic offset that distorts the weight-to-code mapping. With the optimization of the CDAC reference voltages, the full 6-bit code space (0–63) is exercised across the 4-bit weight range. As shown in Fig. 11(b) after post-processing, the calibrated response eliminates the offset present in the uncalibrated case and restores a near-ideal linear mapping. Each weight increment corresponds to an average step of $\sim 4$ ADC codes, thereby preserving monotonicity and maximizing quantization efficiency. Fig. 12 shows the variation in output voltage and current when 128 rows are activated simultaneously for all 1’s in Monte Carlo simulations. These results highlight that reference tuning and calibration are essential to fully exploit the ADC’s dynamic range when interfaced with the analog outputs of the proposed 6T-2R array.
@@ V-D Throughput and Energy Efficiency
The proposed 6T-2R subarray consists of $128\times 512$ bit-cells and follows the input-activation and weight mapping strategy described in the previous section, with both input activation and weight precision set to 4 bits. The system latency is dominated by the SAR ADC, which operates at 50 MHz and incurs a delay of 160 ns. Due to the bit-serial computation approach, the latency for performing PIM using $R_{LEFT}$ is 640 ns, and the same applies for $R_{RIGHT}$. After completing two PIM computations, the normalized throughput achieved is 0.4 TOPS at 0.8 V for maximum array utilization. The overall system energy is primarily dominated by the 6T-2R array, which accounts for approximately 60% of the total energy consumption, followed by the ADC and the WCC block which operates at 1 V. The normalized energy efficiency of the 6T-2R array is calculated to be 452.34 TOPS/W. In terms of area, the total macro size is approximately 0.1 mm2, with the ADC occupying nearly 70% of the area. The resulting normalized compute density is estimated to be 4.37 TOPS/mm2.
@@
Parameter TCASII’24[TCAS2] ISSCC’23[ISSCC23] ISSCC’22[isscc22] TCASI’23[chen2023samba] TCASI’23[song202328] JSSCC’24[nverma] This Work
Technology 180nm CMOS 28nm FDSOI 22nm FDSOI 65nm CMOS 28nm FDSOI 22nm FDSOI 22nm FDSOI
Array Size 8Kb 16Kb 256Kb 101Kb 16Kb 256Kb 64Kb
PIM Domain Time Charge Current Charge Charge Current Current
Memory Type 6T SRAM + 9T 10T1C SRAM 1T1R RRAM 10T1C SRAM 6T SRAM 1T1R MRAM 6T-2R SRAM +RRAM
Cache Data Retention No No No No No No Yes
Dataset CIFAR-10 NA CIFAR-10 CIFAR-10 CIFAR-10 CIFAR-10 CIFAR-10
Accuracy 86.1% NA 91.74% 88.6% 85.07% 90.25% 91.76%
Input/Weight Precision 8/8 8/8 8/8 8/8 4/4 4/4 4/4
@@
Throughput (GOPS) 0.07 7.65 142.2 12.8 12.8 54.3 25.6
Energy Efficiency (TOPS/W) 0.291 16.02 0.96 10.3 16.1 5.26 28.97
@@
Baseline (no ADC nonlinearity or noise) 92.88
ADC nonlinearity only (fine-tuned) 92.00
ADC nonlinearity + noise (fine-tuned) 91.76